### v1

# Argument Summarization and its Evaluation in the Era of Large Language Models

Abstract
Large Language Models (LLMs) have revolutionized various Natural Language Generation (NLG) tasks, including Argument Summarization (ArgSum), a key subfield of Argument Mining (AM). This paper investigates the integration of state-of-the-art LLMs into ArgSum, including for its evaluation. In particular, we propose a novel prompt-based evaluation scheme, and validate it through a novel human benchmark dataset. Our work makes three main contributions: (i) the integration of LLMs into existing ArgSum frameworks, (ii) the development of a new LLM-based ArgSum system, benchmarked against prior methods, and (iii) the introduction of an advanced LLM-based evaluation scheme. We demonstrate that the use of LLMs substantially improves both the generation and evaluation of argument summaries, achieving state-of-the-art results and advancing the field of ArgSum.
1 Introduction
[…]
Thus, there is an urgent need for systematic analysis to understand how LLMs can be effectively utilized for both the generation and evaluation of argument summaries. This includes integrating LLMs into ArgSum frameworks to comprehensively assess their performance and developing suitable prompt-based evaluation schemes.
In this work, we aim to fill this gap by extensively exploring how LLMs can be leveraged for the ArgSum process, both for generating argument summaries and for their evaluation. Our core contributions are: (i) We integrate LLMs into existing ArgSum systems, (ii) we introduce a new LLM-based ArgSum system and benchmark existing work against it, (iii) we provide a new ArgSum evaluation dataset with human evaluation scores, and (iv) we develop a prompt-based ArgSum evaluation scheme and validate it in the context of existing work and our human evaluation.
We show that the use of LLMs substantially improves existing work in both the process of ArgSum and ArgSum evaluation, presenting several state-of-the-art results.
2 Related Work
2.1 Argument Summarization
[…]
Misra et al. (2016), Reimers et al. (2019) and Ajjour et al. (2019) treat the task of summarizing arguments as a clustering problem without providing easy-to-understand textual summaries. Wang and Ling (2016) frame ArgSum as claim generation, where a collection of argumentative sentences is summarized by generating a one-sentence abstractive summary that addresses the shared opinion of the inputs. Schiller et al. (2021) present an aspect-controlled argument generation model that enables an abstractive summatization of arguments.
Our understanding of ArgSum is in line with Key Point Analysis (KPA), introduced by Bar-Haim et al. (2020a, b), and is displayed in Figure 1. They aim to create an extractive summary of the most prominent key points from a potentially large collection of arguments on a given debate topic and stance. Then, each source argument is classified according to the most suitable key point. Alshomary et al. (2021) perform the key point extraction by utilizing a variant of PageRank (Page et al., 1998). Li et al. (2023) extend KPA with a clustering-based and abstractive approach, using grouped arguments as input for a generation model to create key points.
Khosravani et al. (2024) introduce a clustering-based and extractive approach, selecting the most representative argument within each cluster as a key point, determined by a supervised scoring model.
Figure 1: General procedure of ArgSum, where a set of $K$ arguments on a certain debate topic and stance (example taken from Friedman et al. (2021)) is transformed to a set of $N$ argument summaries along with their respective importance (indicated by the size of the green dots). It is expected that $K\gg N$ applies.
2.2 Evaluating NLG Systems
While automatic evaluation metrics such as BLEU Papineni et al. (2002) and ROUGE Lin (2004) correlate poorly with human judgments Novikova et al. (2017), pre-trained transformer-based language models provide a more nuanced assessment of the performance of NLG systems Celikyilmaz et al. (2021). BERTScore Zhang et al. (2020) and MoverScore Zhao et al. (2019) are reference-based metrics that leverage pre-trained embeddings obtained from BERT-based models. While BERTScore is based on the computation of cosine-similarities between the hypothesis and the reference, MoverScore determines an evaluation score by computing the Word Mover’s Distance Kusner et al. (2015) between both. BARTScore Yuan et al. (2021) is based on the pre-trained sequence-to-sequence model BART and treats the evaluation task as a problem of text generation. BLEURT Sellam et al. (2020) is reference-based and consists of a BERT-based model that is fine-tuned to predict human ratings in such a way that the metric is robust to quality shifts. MENLI Chen and Eger (2023) frames the evaluation task as a problem of Natural Language Inference (NLI), showing improved robustness. The most recent approaches to evaluation are LLM-based metrics, which can be leveraged in various ways: by comparing embeddings in terms of their cosine similarity Es et al. (2024), by determining the sequence probability of the hypothesis given the respective source/reference Fu et al. (2024), by utilizing suitable prompting strategies Kocmi and Federmann (2023); Liu et al. (2023); Fernandes et al. (2023); Leiter and Eger (2024); Larionov and Eger (2024), or by applying task-specific fine-tuning Wang et al. (2024); Xu et al. (2023); Zhu et al. (2023). Some works show promising zero-shot results that are on-par with human-judgement Leiter et al. (2023); Chang et al. (2024).
In this work, we leverage LLMs to evaluate ArgSum systems, which is different from evaluation of classical text generation systems, requiring different dimensions of evaluation (e.g., redundancy and coverage) and different mechanisms (e.g., ArgSum requires to compare $m$ reference summaries to $n$ generated summaries). To this end, we apply an LLM-based prompting approach and compare it with two existing ArgSum evaluation frameworks.
2.3 Argument Summarization and Evaluation with LLMs
Among the works that use LLMs for argument summarization or evaluation, Ziegenbein et al. (2024) use snippet generation and neutralization (a mix of extractive and LLM prompting with reinforcement learning) for ArgSum in the context of argument search. They evaluate their approach automatically and manually, but do not apply LLMs for the evaluation. While their ArgSum task is different from ours, they also did not assess state-of-the-art LLMs like GPT-4o. Li et al. (2024) apply LLMs to argumentative essay summarization. They test a variety of state-of-the-art LLMs to generate reference summaries which are evaluated by humans. Their own summarization system, however, relies on smaller, instruction fine-tuned models. Different to our work, for the ArgSum evaluation, they only apply standard metrics like ROUGE.
Our work is the first systematic study on strategies for integrating LLMs with existing approaches for ArgSum and ArgSum evaluation.
3 Experimental Setup
3.1 Terminology
Most previous work on ArgSum can be categorized as either classification-based or clustering-based systems. Classification-based systems first generate a set of argument summaries based on all available source arguments. In a second step, they match each source argument to the most appropriate summary. Clustering-based systems first group all source arguments according to their similarity. Then, they generate a summary of the arguments for each cluster. In this work, we augment ArgSum systems of both types with LLMs. Details of those systems and how we integrate LLMs are specified in §3.2 and §3.3.
The systems we assess use two types of tools to perform ArgSum. While Quality Scorers assess the quality of an argument, Match Scorers determine how well an argument and a summary match. Both are realized by transformer-based language models that take task-specific textual inputs and output a respective score. The ArgSum systems considered in this work utilize Quality Scorers that are fine-tuned on the IBM-ArgQ-Rank-30kArgs (ArgQ) dataset by Gretz et al. (2020). The corresponding Match Scorers are fine-tuned on the ArgKP-2021 (ArgKP21) dataset by Friedman et al. (2021). Details on the required model fine-tuning for the ArgSum systems discussed in §3.2 and §3.3 are collected in Appendix A.
3.2 Classification-based Systems
We consider two classification-based ArgSum systems, which performed best in the Key Point Generation Track at the 2021 Key Point Analysis Shared Task (Friedman et al., 2021).
BarH
To determine a set of potential argument summaries, referred to as candidates, BarH (Bar-Haim et al., 2020b) scores the source arguments with a Quality Scorer and selects those exceeding a threshold $t_{q}$. Subsequently, BarH applies a Match Scorer to match the remaining source arguments to the best fitting candidates. After ranking the candidates according to their number of matches, BarH minimizes redundancy by removing candidates whose match score with a higher-ranked candidate exceeds a threshold $t_{m}$. The remaining candidates are understood as the final argument summaries.
SMatchToPr
To identify argument summary candidates, SMatchToPr (Alshomary et al., 2021) uses a variant of PageRank (Page et al., 1998). To this end, candidates are understood as nodes in an undirected graph, utilizing the match scores between each candidate pair as edge weights. Only nodes with edge weights above a threshold $t_{n}$ are connected. Based on the resulting graph, an importance score is calculated for each candidate. Then, SMatchToPr minimizes redundancy by removing candidates whose match score with a higher-ranked candidate exceeds a threshold $t_{m}$. This results in the final set of argument summaries.
LLM Integration
Given a set of arguments on a certain debate topic and stance, we apply a zero-shot prompting approach to instruct an LLM to generate either a set of candidates or argument summaries (see Appendix B.1). The resulting candidates or argument summaries are then further processed as usual in both BarH and SMatchToPr.
3.3 Clustering-based Systems
We consider an approach from Li et al. (2023) which demonstrated comparable performance to BarH and SMatchToPr. Further, we propose a new ArgSum approach that utilizes a Match Scorer for argument clustering.
USKPM
For clustering arguments, USKPM (Li et al., 2023) utilizes the BERTopic framework (Grootendorst, 2022), which involves three steps. First, contextualized sentence embeddings of the arguments are created via SBERT (Reimers and Gurevych, 2019). Second, UMAP (McInnes et al., 2018) is applied to reduce the embeddings’ dimensionality. Third, the clustering of the reduced embeddings is performed by HDBSCAN (McInnes et al., 2017). Instances included in clusters with a size smaller than $c$ are considered as unclustered. Since Li et al. (2023) state that it is reasonable to maximize the number of clustered arguments in order to increase the representativeness of the argument summaries to be generated, Iterative Clustering (IC) is proposed. IC is about incrementally assigning unclustered arguments to the most similar cluster in terms of cosine similarity.
Then, USKPM uses the instruction-tuned FLAN-T5 (Chung et al., 2022) to summarize the argument clusters, where the model input is formatted as follows: “summarize: {Stance} {Topic} {List of Arguments in Cluster}”.
MCArgSum
Our own approach, MCArgSum (Match Clustering based ArgSum), combines the use of a Match Scorer for argument clustering with an LLM-based cluster summarization. It is inspired by the redundancy reduction among candidates within BarH, where a Match Scorer is utilized to identify candidates addressing the same key point. We demonstrate that a Match Scorer can also be effectively used to group arguments addressing the same main statement. While the key idea of using a Match Scorer to group arguments is also proposed by Khosravani et al. (2024), our ArgSum system additional provides an abstractive summarization of argument clusters by incorporating an LLM.
Our ArgSum system utilizes Agglomerative Hierarchical Clustering (Day and Edelsbrunner, 1984) with the average linkage criterion in reference to Reimers et al. (2019) and a Match Scorer as pairwise similarity metric. To this end, we use the SBERT (Reimers and Gurevych, 2019) model all-mpnet-base-v2, fine-tuned on ArgKP21. While the threshold $m$ determines the minimum match score required between two clusters to be merged, instances included in clusters with a size smaller than $c$ are considered as unclustered.
To generate cluster summaries, our model uses LLM prompting in a zero-shot setting. We integrate a prompting strategy that summarizes all argument clusters simultaneously (global optimization). Details are given in Appendix B.2. After summarization, a post-processing step automatically extracts the argument summaries in the desired format.
3.4 Evaluation
Here, we describe the approaches used to evaluate ArgSum systems. These metrics are both set-based and reference-based, meaning a set of candidate summaries is compared to a set of reference summaries.
In accordance with previous work on generating argument summaries, we assess the two evaluation dimensions of coverage and redundancy. Coverage refers to the extent to which a set of argument summaries captures the central talking points of a debate. Redundancy is concerned with the extent of content overlap between the individual argument summaries (Bar-Haim et al., 2020b; Alshomary et al., 2021; Friedman et al., 2021; Li et al., 2023; Khosravani et al., 2024). Both criteria are closely related in order to assess the overall quality of a set of argument summaries, as high coverage can be achieved by generating many redundant argument summaries (Friedman et al., 2021).
Soft-Score
Li et al. (2023) introduce three evaluation scores: Soft-Precision (sP), Soft-Recall (sR) and Soft-F1 (sF1). While sP finds the most suitable reference summary for each candidate summary, sR finds the most suitable candidate summary for each reference summary. To compare references and candidates, Li et al. (2023) utilize a semantic similarity function. The final evaluation scores in terms of sP and sR are obtained by averaging the similarity scores of the respective best matches of references and candidates. Finally, the sF1 is the harmonic mean of sP and sR. Formally:
$sP=\frac{1}{n}\cdot\sum_{a_{i}\in A}\max_{\beta_{j}\in B}f(a_{i},\beta_{j})$ (1)
$sR=\frac{1}{m}\cdot\sum_{\beta_{j}\in B}\max_{a_{i}\in A}f(a_{i},\beta_{j})$ (2)
where $f$ is a function evaluating the semantic similarity between two summaries; $A$ and $B$ are the sets of candidate and reference summaries, with $n$ and $m$ being their respective sizes. As similarity function, Li et al. (2023) suggest the use of BLEURT (Sellam et al., 2020) and BARTScore.
Coverage-Score (CS)
The Coverage-Score (CS) (Khosravani et al., 2024) assesses the coverage of a set of candidate summaries, which is defined as the proportion of reference summaries covered by them. Each possible pair of candidates and references is scored by a Match Scorer and classified as matching or non-matching. The former corresponds to the case in which the respective match score is above a certain threshold $t_{M}$. Finally, the CS is derived as the proportion of references with at least one matching candidate. Formally:
${CS=\frac{1}{m}\sum_{\beta_{j}\in B}{\mathds{1}}\left[\sum_{a_{i}\in A}{\mathds{1}}\left[\textit{match}(a_{i},\beta_{j})>t_{M}\right]\geq 1\right]}$ (3)
where match indicates the match score of two summaries; $A$ and $B$ are the sets of candidate and reference summaries, $m$ is the size of $B$ and $t$ is the matching threshold. Khosravani et al. (2024) suggest the use of the Match Scorer inherent in BarH. A recommended threshold $t_{M}$ is not provided.
LLM-based
We introduce two one-shot prompting strategies for assessing ArgSum systems, focusing on the dimensions of coverage and redundancy. (i) We address coverage by instructing an LLM to count the number of reference summaries covered by a set of candidate summaries. Dividing this count of covered references by the total number of references results in an LLM-based coverage score. (ii) To assess redundancy, we instruct an LLM to count the number of unique main statements within a set of candidate summaries. The resulting uniqueness count is limited to the total number of candidates and a uniqueness score is derived by dividing the uniqueness count by the total number of candidates. Subsequently, we derive an LLM-based redundancy score as the complementary uniqueness score $(1-\textit{uniqueness})$. The final LLM-based coverage and redunancy scores for a certain set of candidate summaries is obtained by averaging the results of 10 evaluation runs.
Human Evaluation
To verify the reliability of the automatic evaluation metrics, we conduct a human evaluation of 126 generated argument summaries obtained from the ArgSum systems described in §3.2 and §3.3. We characterize a suitable set of argument summaries as consisting of succinct, non-redundant summaries that cover the main statements shared across the source arguments with adequate granularity. Thus, we assess the dimensions of coverage and redundancy, as introduced above.
The judgments are carried out by four experienced annotators with excellent knowledge within the field of NLP, especially argumentation. Initially, the four annotators are introduced to the task of ArgSum and provided with a description of the evaluation task. Guidelines can be found in Appendix C. The annotators are presented with a set of argument summaries (generated by the ArgSum systems discussed in Sections §3.2 and §3.3 and based on the arguments contained in ArgKP21) and the corresponding set of reference summaries. To assess coverage, they are asked to count the number of references that are covered by the set of generated summaries. The respective coverage score is the proportion of covered references out of the total number of references. We then ask the annotators to count the number of unique main statements within the set of generated summaries (permitted count is limited to the total number of generated summaries). Based on this, we derive a uniqueness score (ranging from zero to one) as the number of unique main statements divided by the total number of generated summaries. The redundancy score is the complementary uniqueness score.
In order to determine the inter-rater reliability, we average the Pearson correlation coefficients between each pair of the four annotators’ scores for both dimensions. We report an average correlation of 0.697 for coverage and 0.722 for redundancy, indicating that the annotations are reliable. Pairwise correlations between annotators are shown in Figure 3 in the appendix. [^2] For annotator 1 (A1), the judgments for both stances of the third topic are missing, whereas the others (A2-A4) evaluated all three topics.
4 Results
In this section, we present the correlation of automatic metrics with human judgments in §4.1 and the evaluation of ArgSum systems in §4.2. Details on the experimental conditions, including data preprocessing, modifications to the ArgSum systems, hyperparameter settings, and hardware, are provided in Appendix D
Data
While ArgKP21 is used to train the Match Scorers utilized by BarH, SMatchToPr and MCArgSum, we use its test set to generate argument summaries in §4.1 and §4.2.
This dataset consists of 27,519 pairs of arguments and key points, each labeled with a binary value that indicates whether the corresponding argument and key point are matching (1) or non-matching (0). While the pairs of arguments and key points cover 28 topics, each with supporting (1) and opposing (-1) stance, the dateset includes a train set of 24 topics, a development set of 4 topics and a test set of 3 topics.
We also consider the Debate dataset (Hasan and Ng, 2014) as a second independent evaluation data set in §4.2. Debate includes 3,228 argumentative text sequences filtered from posts on four different topics in an online debate forum. The text sequences are labeled with their reason within the respective topic and whether they are supporting (1) or opposing (-1). We consider the argumentative text sequences as arguments and the reasons as argument summaries. In contrast to ArgKP21, the dataset exclusively contains matching pairs.
Exemplary data points for ArgKP21 and Debate are presented in Table E and Table E in the appendix, respectively.
Representative LLM
As representative LLM, we utilize GPT-4o (gpt-4o-2024-08-06) for ArgSum, as well as GPT-4o-mini (gpt-4o-mini-2024-07-18) for evaluation, both provided by OpenAI’s API service. [^3] https://platform.openai.com/docs/models/gpt-4o-mini GPT-4o offers more qualitative responses for a suitable generation of argument summaries, while GPT-4o-mini was chosen for evaluation as it offers fast response times and is a cost-effective model version for the more quantity-based evaluation approach. We integrated GPT-4o into the ArgSum systems as described in §3.2 and §3.3. GPT-4o-mini was integrated into the LLM-based evaluation metric as discussed in §3.4. Nevertheless, we continue to use the general term “LLM” to indicate that GPT-4o can be replaced with any other generative LLM.
4.1 Reliability of automatic metrics
To measure the quality of diverse automatic metrics, we correlate them to our human assessment of 126 argument summaries, see §3.4, where we average the four human assessments per instance, focusing on coverage as annotated by humans.
We consider two ways of computing correlations. (i) We calculate correlations across all topics and stances simultaneously. (ii) We calculate correlations within topics and stances and average the results. For the latter scenario, we also report the standard deviations, indicating the variability of reliability.
Soft-Score
We apply the Soft-Score, explained in §3.4, with the following automatic metrics as similarity function $f$: (1) ROUGE 1, (2) BERTScore F1 (Zhang et al., 2020), (3) MoverScore (Zhao et al., 2019), (4) BARTScore (Yuan et al., 2021), (5) BLEURT (Sellam et al., 2020), (6) MENLI (Chen and Eger, 2023). Table E in the appendix shows the results.
First, we note that sP does not intuitively correspond to the annotation dimensions of coverage or redundancy in our human annotation — sP could be interpreted as the fraction of candidate summaries covered by the reference summaries, but not vice versa. Thus, it comes as no surprise that the correlation between sP and coverage is close to zero across all settings. The sR, which better matches the definition of coverage, performs clearly better, even if no strong correlations are observed. Across topics and stances, MENLI performs best (0.265) followed by BERTScore-F1 (0.254). The scenario within topics and stances generally yields better correlation results for the sR. While BERTScore-F1 exhibits the highest correlation at 0.402, MENLI (0.372) also achieves a moderate positive correlation with the human coverage scores. It is notable that BLEURT and BARTScore, suggested by Li et al. (2023), achieve the poorest results among all considered similarity functions. [^4] We rescaled BARTScore according to Li et al. (2023) in order to obtain positive scores in the range from zero to one.
{tblr}
[…]
Across Within Runtime (s) Across Within Runtime (s)
0.20 0.736 0.756 $\pm$ 0.100 495.1 0.798 0.697 $\pm$ 0.226 1305.3
[dotted] 0.30 0.725 0.747 $\pm$ 0.133 467.7 0.789 0.752 $\pm$ 0.096 1651.1
[dotted] 0.40 0.746 0.771 $\pm$ 0.112 529.3 0.817 0.758 $\pm$ 0.122 1515.4
[dotted] 0.50 0.742 0.757 $\pm$ 0.127 512.0 0.812 0.724 $\pm$ 0.210 1446.3
[dotted] 0.60 0.741 0.762 $\pm$ 0.122 629.7 0.837 0.795 $\pm$ 0.088 1359.9
[dotted] 0.70 0.755 0.789 $\pm$ 0.103 644.3 0.830 0.782 $\pm$ 0.112 1425.6
[dotted] 0.80 0.729 0.755 $\pm$ 0.108 612.4 0.828 0.762 $\pm$ 0.111 1431.1
[dotted] 0.90 0.754 0.782 $\pm$ 0.131 676.4 0.843 0.784 $\pm$ 0.109 1651.0
[dotted] 1.00 0.767 0.803 $\pm$ 0.115 845.5 0.852 0.824 $\pm$ 0.055 1649.1
Table 1: Pearson correlation between the LLM-based coverage and redundancy scores and the respective averaged human scores for different temperatures, along with the evaluation runtime, on ArgKP21. For the scenario within topics and stances, standard deviations are indicated alongside the correlation values.
Coverage Score
To examine the correlation of the CS with the averaged human coverage scores, we consider the Match Scorers of BarH, as proposed by Khosravani et al. (2024), as well as those of SMatchToPr and MCArgSum. Furthermore, we apply various values for the threshold $t_{M}$, which determines the match score for which an individual reference summary is understood as covered or not.
As depicted in Table E in the appendix, the CS with BarH’s Match Scorer reaches a maximum correlation of 0.489 across and 0.698 within topics and stances. For the scenario across topics and stances, SMatchToPr performs even better and achieves a maximum correlation of 0.541. Within topics and stances, SMatchToPr reaches a maximum correlation of 0.6. The Match Scorer included in MCArgSum yields comparatively worse results, achieving a maximum correlation of 0.449 within and 0.551 across topics and stances. Regarding the matching threshold $t_{M}$, BarH’s Match Scorer performs very stably across the considered parameter range, whereas this is not the case for both other variants.
To summarize, the CS provides considerably stronger correlations for the dimension of coverage compared to the Soft-Score.
LLM-based metric
The LLM-based metrics for coverage and redundancy, described in §3.4, are examined regarding their respective criterion. Here, we investigate different values for the temperature, a parameter controlling the creativity or randomness in LLM-based text generation (Peeperkorn et al., 2024). The results are collected in Table 4.1.
The LLM-based score for coverage achieves a maximum correlation of 0.767 across and 0.803 within topics and stances. Consequently, it performs better than the Soft-Score and CS in all scenarios. The LLM-based metric for redundancy reaches also high correlations with a maximum value of 0.852 across and 0.824 within topics and stances. Thus, we exclusively use the LLM-based evaluation metrics to assess the argument summarization capability of ArgSum systems in §4.2.2.
4.2 System evaluation
In this section, we evaluate the argument separation capability of clustering-based ArgSum systems in §4.2.1 and then assess the ArgSum systems’ argument summarization capability in §4.2.2.
4.2.1 Argument Separation Capability
We first evaluate the capability of clustering-based ArgSum systems (USKPM, MCArgSum, as described in §3.2, along with two MCArgSum variants using the Match Scorers of BarH and SMatchToPr) to separate arguments. To determine the argument separation capability of ArgSum systems, we use the Adjusted Rand Index (ARI) (Warrens and van der Hoef, 2022), which measures the agreement between two data clusterings while accounting for the possibility of random clustering. The ARI ranges from -1 to 1, where lower values indicate poor separation, higher values indicate good separation, and values near zero suggest random clustering.
Since both processed test datasets consist exclusively of arguments with exactly one matching summary, we treat the sets of arguments associated with each unique argument summary as reference clusters within topics and stances. The argument clusterings provided by the considered ArgSum systems substantially depend on the hyperparameters used. Therefore, we conduct several clustering runs with different parameter settings and select the best runs within topics and stances for each ArgSum system. The parameter settings taken into account are described in Appendix D.3.
As the considered ArgSum systems do not force each instance to be clustered, it is possible that a run in which only a small proportion of arguments is clustered achieves a high ARI. To take this into account, we consider two scenarios to determine the best clustering runs. First, we average the highest ARIs of each topic and stance, provided that the respective proportion of clustered arguments is above 70%. Since this scenario excludes unclustered arguments, we refer to it as excluding noise. Second, we include noise by considering unclustered arguments as a separate cluster and averaging the highest ARI across topics and stances.
The results for ArgKP21 and Debate are presented in Table E and Table E as well as Figure 4 and Figure 5 in the appendix. Depending on the Match Scorer, MCArgSum performs 0.312-0.452 ARI points better than USKPM on ArgKP21 and 0.07-0.113 ARI points better on Debate when considering the scenarios including and excluding noise. Only when comparing MCArgSum (SMatchToPr) in the scenario including noise to USKPM with IC on Debate does USKPM perform slightly better by 0.07 ARI points. Although MCArgSum (BarH) and MCArgSum (SMatschToPr) yield the best results on ArgKP21, and MCArgSum (BarH) outperforms MCArgSum as presented in §3.2 on Debate, our variant offers significant advantages in terms of time efficiency. While the best performing MCArgSum (BarH), for example, requires an average runtime of 201.9 seconds per topic and stance on Debate, our MCArgSum variant requires an average of only 1.9 seconds.
To summarize, we demonstrate that our Match Scorer-based clustering approach can separate arguments substantially better than the existing argument clustering approach included in USKPM, and that the variant of MCArgSum as described in §3.2 comes with huge advantages in terms of time effeciency.
4.2.2 Argument Summarization Capability
BarHBarH+candBarH+summSMtPRSMtPR+candSMtPR+summUSKPMMCArgSum000.50.50.50.511110.820.820.820.820.90.90.90.90.810.810.810.810.910.910.910.910.910.910.910.910.80.80.80.80.820.820.820.820.840.840.840.840.090.090.090.090.180.180.180.180.050.050.050.050.240.240.240.240.170.170.170.170.050.050.050.050.250.250.250.250.160.160.160.160.850.850.850.850.880.880.880.880.860.860.860.860.860.860.860.860.880.880.880.880.850.850.850.850.80.80.80.80.840.840.840.84CoverageRedundancyWeighted
BarHBarH+candBarH+summSMtPRSMtPR+candSMtPR+summUSKPMMCArgSum000.50.50.50.511110.760.760.760.760.840.840.840.840.80.80.80.80.780.780.780.780.860.860.860.860.790.790.790.790.810.810.810.810.880.880.880.880.220.220.220.220.150.150.150.150.080.080.080.080.150.150.150.150.120.120.120.120.040.040.040.040.110.110.110.110.110.110.110.110.770.770.770.770.850.850.850.850.840.840.840.840.810.810.810.810.870.870.870.870.840.840.840.840.830.830.830.830.890.890.890.89CoverageRedundancyWeighted
Figure 2: Coverage and redundancy scores as well as Weighted Scores for ArgKP21 (top) and Debate (bottom). For BarH and SMatchToPr (abbreviated as SMtPR), the variants with LLM-based candidates and summaries are indicated by +cand and +summ, respectively.
Having identified the LLM-based evaluation metrics as the most reliable among those considered for both dimensions of coverage and redundancy, this section addresses their application in order to evaluate the ArgSum systems. In our investigations, we make use of a weighted evaluation score assessing both coverage and redundancy simultaneously. The Weighted Score $ws$ for a certain set of argument summaries is defined as follows:
$ws=\alpha\cdot c+(1-\alpha)\cdot(1-r)$ (4)
where $c$ indicates the LLM-based coverage score and $r$ indicates the LLM-based redundancy score. The weighting factor $\alpha$ is defined to be in the range [0, 1] and can be used to bias the Weighted Score either towards the coverage score or the redundancy score. For our investigations, we set the weighting factor to $2/3$, as we consider coverage to be more important than redundancy. We generate several argument summaries using various hyperparameter settings (see Appendix D.3) and report the best setting in terms of the Weighted Score for each ArgSum system. Since ArgSum is performed per topic and stance, the final evaluation score for each ArgSum system results as the average of the highest Weighted Scores within topics and stances. For simplicity, we refer to the averaged highest Weighted Score as the Weighted Score and the averaged coverage and redundancy score as the coverage and redundancy score, respectively.
The results of evaluating the ArgSum systems with regard to their generated arguments summaries for ArgKP21 are depicted in Figure 2 (top). For both classification-based systems, the integration of an LLM leads to improved performance. The variant of BarH without LLM integration has a slightly lower Weighted Score (0.85) compared to the LLM-based summary generation results (0.86). The variant of BarH with LLM-based argument candidate generation performs best and achieves a Weighted Score of 0.88. Considering SMatchToPr, the variant with LLM-based argument candidate generation (0.88) also outperforms those with LLM-based summary generation (0.85) and without LLM integration (0.86). It is noteworthy that SMatchToPr without the integration of an LLM achieves the highest coverage score (0.91) among all ArgSum systems, but also comes with second highest redundancy (0.24). With regard to the clustering-based ArgSum systems, MCArgSum achieves a Weighted Score of 0.84 and outperforms USKPM (0.80). Nevertheless, any classification-based system delivers better results.
Considering the results for Debate, shown in Figure 2 (bottom), the integration of LLMs comes with advantages in all cases. As with ArgKP21, the variants of integrating LLMs for argument candidate generation generally performs best when looking at classification-based ArgSum systems. However, the clustering-based MCArgSum provides even better results: It achieves a Weighted Score of 0.89, mostly as a result of its high coverage score (0.88).
The integration of LLMs results in considerable improvements for classification-based as well as clustering-based ArgSum systems. It is notable that LLM-based argument candidate generation in classification-based systems performs best for both datasets. MCArgSum is particularly convincing on Debate and achieves the best result here.
The final choice of an ArgSum system should also depend on the runtime requirements. Here, clustering-based systems are generally faster, with MCArgSum showing the best performance among all LLM-based ArgSum systems for both datasets. It required on average 3.779 seconds per topic and stance for ArgKP21 and 7.375 seconds for Debate (cf. hardware specifications in Appendix D.4).
5 Conclusion
Our proposed LLM-based ArgSum systems and metrics achieve state-of-the-art performance across the two datasets considered. MCArgSum, our newly proposed LLM-based ArgSum system, achieves highest performance on the Debate dataset and has a runtime advantage against all other systems considered. The LLM-based ArgSum evaluation scores we propose show very high correlation with human judgements and thus set a very reliable evaluation framework where reference summaries are available.
A few open questions and tasks remain: we did not consider open-source LLMs like Llama 3.X or Deepseek R1. While we do not expect substantial improvements to GPT-4o, it might be interesting to better understand the influence of the LLM on the prompting strategies relevant for the ArgSum systems and (even more) evaluation. Furthermore, we leave the application of reference-free evaluation strategies to future work.
Limitations
Both model’s (GPT-4o (gpt-4o-2024-08-06), GPT-4o-mini (gpt-4o-mini-2024-07-18)) training data includes information up to October 2023. ArgKP21, published in November 2021 (Friedman et al., 2021) and Debate, which dates back to 2014 (Hasan and Ng, 2014) could have been used in training. However, similar limitations of potential data contamination are faced in many other recent problem settings as well; due to a lack of suitable ArgSum datasets, this issue is hard to avoid. We also point out that this work introduces a new evaluation benchmark for ArgSum systems, which could not have been seen by our employed LLMs.
Ethical Considerations
ArgSum systems could yield unreliable, factually incorrect or even maliciously misleading summaries of the underlying source arguments. Thus, recipients of the summarized arguments must interpret these with care.
Appendix A Details on Model Fine-tuning
BarH and SMatchToPr
We fine-tuned the Match Scorers and Quality Scorers in BarH and SMatchToPr according to Bar-Haim et al. (2020b) and Alshomary et al. (2021), respectively. It is important to note that Bar-Haim et al. (2020b) do not specify which of the two quality scores (MACE-P and WA) in ArgQ should be used for training the Quality Scorer. Additionally, it is unclear whether a model with or without a pooling layer was used. Since the model without pooling layer and fine-tuned on MACE-P performs best in preliminary investigations, we applied it in BarH.
USKPM
The fine-tuning of FLAN-T5 in USKPM was conducted as proposed by Li et al. (2023), though no specific learning rate was provided. Based on our observations, a learning rate of 4e-4 worked well and was therefore used for fine-tuning the model.
MCArgSum
As Match Scorer, MCArgSum uses the SBERT model “all-mpnet-base-v2” fine-tuned on ArgKP21. The fine-tuning is conducted over 10 epochs with a learning rate of 5e-6 and contrastive loss. The best performing model on the development set was selected as final model.
Appendix B LLM Prompting
LLM prompting can be divided into a system message and a user message. The system message guides the LLM on its general behavior, while the user message specifies the exact task. We also utilize the system message to introduce the task at hand and to describe the desired appearance of the argument summaries.
B.1 Classification-based Systems
The proposed prompting strategy instructs the LLM to generate either candidates or argument summaries. In both cases, the prompt is divided into a system message and a user message. The following prompt template is applied for both generating candidates and argument summaries, but used differently. For generating candidates, we instruct the LLM to produce a large number of key points (12 to 20). In contrast, for argument summaries, we request fewer key points (4 to 8) and apply the optional user message to minimize redundancy. A description of the parameters and placeholders contained in the prompt template is given below.
System Message
You are a professional debater and you can express yourself succinctly. If you are given a corpus of arguments on a certain debate topic and stance, you find {num_kps} appropriate salient single sentences, called key points, summarizing most of the arguments and providing a textual and quantitative view of the data. A key point can be seen as a meta argument why one is for or against a certain topic. Make sure that the generated key points summarize the majority of the arguments contained in the corpus. A key point should not exceed a length of {kp_token_length} tokens. Here are two examples of good key points: “School uniform reduces bullying” is an opposing key point on the topic “We should abandon the use of school uniform" and "Guns lead to accidental deaths” is a supporting key point on the topic "We should abolish the right to keep and bear arms".
User Message
Please generate {num_kps} short (maximal length of {kp_token_length} tokens), salient and high quality {stance} key points on the topic “{topic}” so that they capture the main statements that are shared between most of the arguments based on the following corpus of arguments: {arguments}.
Optional User Message for generating argument summaries
You should only generate as many key points as necessary to summarize the arguments contained in the corpus. This means you should preferably generate fewer key points than the maximum permitted number of {max_num_kps} key points instead of generating overlapping key points in terms of content.
Parameters/Placeholders
  • num_kps: Number of key points (can be a fixed value or a range of values)
  • kp_token_length: Maximum permitted number of tokens for key points
  • stance: Stance of arguments (supporting or opposing)
  • topic: Topic of arguments
  • arguments: List of arguments
  • max_num_kps: Maximum permitted number of key points
B.2 Clustering-based Systems
The prompting for LLM-based Cluster Summarization is divided into a system message and a user message. A description of the parameters and placeholders contained in the prompt template is given below.
System Message
You are a professional debater and you can express yourself succinctly. If you are given a cluster of similar arguments on a certain debate topic and stance, you find a single appropriate salient sentences, called key point, capturing the main statement that is shared between most of the clustered arguments and providing a textual and quantitative view of the data. A key point can be seen as a meta argument why one is for or against a certain topic. Since argument clusters are not perfect, they may contain arguments that do not actually belong together. Therefore, make sure that a generated key point summarizes the majority of the arguments contained in the cluster. A key point should not exceed a length of {kp_token_length} tokens. Here are two examples of good key points: “School uniform reduces bullying” is an opposing key point on the topic “We should abandon the use of school uniform" and “Guns lead to accidental deaths” is a supporting key point on the topic “We should abolish the right to keep and bear arms”.
User Message
Please generate a single short (maximal length of {kp_token_length} tokens), salient and high quality {stance} key point on the topic “{topic}” so that it captures the main statement that is shared among most of the clustered arguments for each of the following {num_clusters} clusters of similar arguments: {clusters}. Since argument clusters are not perfect, they may contain arguments that do not actually belong together. Therefore, make sure that each generated key point summarizes the majority of the arguments contained in the respective cluster. In addition, ensure that the generated key points do not overlap in terms of content. Do not deliver an explanation why you generated the key points or any other information. Only return the cluster ids and corresponding individual key points.
Parameters/Placeholders
  • kp_token_length: Maximum permitted number of tokens for key points
  • stance: Stance of arguments (supporting or opposing)
  • topic: Topic of arguments
  • arguments: List of arguments
  • num_clusters: Number of clusters
  • clusters: List of argument clusters, where each cluster consists of a cluster id and a list of the corresponding arguments
B.3 LLM-based Evaluation
Appendix C Human Evaluation
C.1 Introduction to ArgSum
A debate on a certain topic can be conducted using a variety of arguments for each side of the debate. Although some of these arguments refer to the same main statement, they can be formulated very differently. While the number of possible arguments seems to be almost infinite due to the possibility of different formulations, the number of possible main statements within a debate is limited.
Argument summarization is about summarizing a relatively large set of arguments on a certain debate topic and stance by generating a small set of argument summaries, each expressing one distinct main statement contained in the set of arguments. In addition, each argument is matched to the generated summary that conveys its main statement the best. Following is a simple example:
Topic:
We should abandon the use of school uniform
Stance:
Opposing
Set of Arguments:
  1. 1.
    School uniforms keep everyone looking the same and prevent bullying
  2. 2.
    School uniforms can help parents save money on outfit
  3. 3.
    School uniforms help stop bullying because when people are similarly dressed, nobody is made to feel inferior
  4. 4.
    It is cheaper for parents to buy school uniforms, which is helpful to parents that are struggling financially
  5. 5.
    School uniforms are substantially more affordable
Set of Summaries:
  1. 1.
    School uniforms reduce bullying
  2. 2.
    School uniforms save costs
Argument Summary Matches:
The matches are highlighted by the colored markings:
  • Arguments 1 and 3 are matched to summary 1
  • Arguments 2, 4 and 5 are matched to summary 2
C.2 Description of the Evaluation Task
This task is about determining how well a set of generated argument summaries serves as a summary of possible arguments on a certain debate topic and stance.
For this purpose, you are given a set of generated summaries and a set of reference summaries as well as the corresponding debate topic and stance. You have to carry out the following two instructions regarding the dimensions of coverage and uniqueness:
  1. 1.
    Coverage: Count the number of reference summaries that are covered by the set of generated summaries.
  2. 2.
    Uniqueness: Count the number of distinct/unique main statements contained in the set of generated summaries.
For both dimensions increments of 0.5 are allowed. In the case of coverage, this applies if a reference summary is only partially covered by the set of generated summaries. For the dimension of redundancy, this applies if there is a distinct main statement in the set of generated summaries that partially overlaps with another. For the case you are not sure, you can answer with -1. Following is an example:
Topic:
Routine child vaccinations should be mandatory
Stance:
Opposing
Set of Reference Summaries:
  1. 1.
    Mandatory vaccination contradicts basic rights
  2. 2.
    Routine child vaccinations are not necessary to keep children healthy
  3. 3.
    Routine child vaccinations, or their side effects, are dangerous
  4. 4.
    The parents and not the state should decide
Set of Generated Summaries:
  1. 1.
    Vaccinations violate free will and personal choice
  2. 2.
    Mandatory vaccines conflict with religious beliefs
  3. 3.
    Parents should have the right to decide
  4. 4.
    Children may suffer harmful effects from vaccines
  5. 5.
    Concerns about vaccine safety and side effects
Coverage:
3 (The second reference summary is not covered.)
Uniqueness:
3.5 (The first and third generated summaries address two different distinct main statements. The fourth and fifth generated summaries refer to the same distinct main statement. The second generated summary partially overlaps with the first one.)
Appendix D Experimental conditions
D.1 Data Preprocessing
To conduct our investigations on the test split of ArgKP21 as well as Debate, we performed two pre-processing steps. First, we remove arguments that do not have exactly one matching argument summary. The reason for this is that we aim to process only those arguments that have a well-defined reference summary. This is because the considered automatic evaluation metrics are reference-based. Including arguments without any reference could result in candidate summaries that are not captured by the references and thus bias the evaluation of ArgSum systems. Beyond that, including arguments with multiple references is not suitable for the evaluation of the argument separation capability addressed in §4.2.1.
Second, we exclude arguments consisting of more than one sentence, as we consider an adequate argument to consist of a single sentence. This is particularly crucial for the argumentative text sequences contained in Debate. For the test split of ArgKP21, the pre-processing reduces the number of arguments from 732 to 428, while for Debate it is reduced from 3180 to 2321. Finally, to decrease the computational effort, we select only 50% of the arguments for each unique argument summary in Debate as our final dataset. This pre-processing step results in 1165 remaining arguments for Debate, while retaining each unique argument summary.
D.2 Modifications to ArgSum Systems
We had to apply three modifications to the ArgSum systems as proposed in §4.2.2. The first concerns the candidate selection in BarH and SMatchToPr. In cases where the proportion of candidates out of all arguments is below a certain threshold $p_{C}$, we fill this gap with the highest quality arguments not yet considered as candidates. In this way, we avoid cases in which no candidates are identified at all, as the Quality Scorer provides low scores across all arguments. Second, when selecting candidates in SMatchToPr, we delete arguments consisting of several sentences instead of separating them. Finally, we use the Quality Scorer included in BarH instead of TextRank for determining the order of arguments in the corresponding input list of Flan-T5 in USKPM.
D.3 Details on Hyperparameters
When applying BarH and SMatchToPr, we used the recommended parameter values from Bar-Haim et al. (2020b) and Alshomary et al. (2021), respectively. In case of USKPM and MCArgSum, we set the minimum cluster size $c$ to 3. The similarity threshold for IC in USKPM was set to zero, meaning that we forced each unclustered argument to be assigned to an existing cluster. In addition, Appendix E includes the varying hyperparameter settings for the argument clustering inherent in USKPM and MCArgSum. For USKPM, we performed the clustering for each possible combination of the depicted parameter values.
D.4 Hardware
We conducted our experiments on a personal computer with an Apple M1 Max chip, which is designed as a system-on-a-chip. It includes a 10-core CPU (8 performance cores and 2 efficiency cores), a 32-core GPU, and a 16-core Neural Engine. The GPU has direct access to the entire main memory of 64GB. The system runs on macOS Sonoma 14.1.2 (64-bit). With the introduction of Metal support for PyTorch on macOS, utilizing the GPU for machine learning tasks has become accessible. [^5] https://pytorch.org/blog/introducing-accelerated-pytorch-training-on-mac This setup was used for both training and inference of PyTorch models.
Appendix E Tables and Figures
{tblr}
colspec = X[c,m,-1]|X[c,m]|X[c,m]|X[c,m], width=1stretch = 0, rowsep = 4pt, Parameter Value Range Steps
[…]
[dotted] Number of neighboring samples used for the manifold approximation of UMAP $[2,5]$ $1$
[dotted] Minimum permitted distance of points in the low dimensional representation of UMAP $[0,0.4]$ $0.2$
[…]
Table 2: Hyperparameter settings of clustering-based ArgSum systems considered in our investigations.
Figure 3: Pairwise Pearson correlation coefficient of the human judgments by the four annotators (A1-A4) for the criteria of coverage and redundancy. The averaged value across annotator pairs is indicated in the parentheses.
{tblr}
colspec = X[l,m,-1]|X[l,m], width=1stretch = 0, rowsep = 4pt, Topic We should abandon the use of school uniform
[dotted] Stance -1
[dotted] Argument school uniforms cut down on bulling and keep everyone the same.
[dotted] Key Point School uniform reduces bullying
[dotted] Label 1
[dotted] Set dev
Table 3: Exemplary data point of ArgKP21.
{tblr}
colspec = X[l,m,wd=21mm]|X[l,m], width=1stretch = 0, rowsep = 4pt, Topic obama
[dotted] Stance -1
[dotted] Argument Where are those outspoken democrats who voted for him because they were told, no promised, that he would END THE WAR?
[dotted] Argument
Summary
Wars are still on
Table 4: Exemplary data point of Debate.
{tblr}
[…]
time (s)
Across Within Across Within Across Within
ROUGE 1 -0.118 -0.072
$\pm$0.194 0.164 0.315
$\pm$0.170 0.027 0.127
$\pm$0.184 0.428
[dotted] BERTSc.
F1 -0.028 0.092
$\pm$0.262 0.254 0.402
$\pm$0.175 0.121 0.240
$\pm$0.242 354.2
[dotted] MoverSc. -0.046 0.044
$\pm$0.227 0.156 0.310
$\pm$0.204 0.069 0.191
$\pm$0.207 55.93
[dotted] BARTSc.
CNN/DM -0.146 -0.305
$\pm$0.164 0.024 -0.011
$\pm$0.283 -0.053 -0.132
$\pm$0.264 84.33
[dotted] BARTSc.
Parabank -0.271 -0.221
$\pm$0.251 -0.012 0.112
$\pm$0.339 -0.132 -0.022
$\pm$0.320 41.09
[dotted] BLEURT -0.209 -0.218
$\pm$0.289 0.033 0.138
$\pm$0.247 -0.091 -0.055
$\pm$0.294 487.3
[dotted] MENLI -0.154 -0.039
$\pm$0.287 0.265 0.372
$\pm$0.260 0.107 0.228
$\pm$0.298 254.6
Table 5: Pearson correlation between the Soft-Score (incl. different similarity functions) and averaged human coverage scores, along with the evaluation runtime, on ArgKP21. For the scenario within topics and stances, standard deviations are indicated below the correlation values.
{tblr}
[…]
Across Within Across Within Across Within
0.40 0.478 0.585 $\pm$0.254 -0.092 -0.157 $\pm$0.037 0.206 -0.057 $\pm$0.223
[dotted] 0.45 0.475 0.605 $\pm$0.248 0.163 -0.074 $\pm$0.187 0.300 -0.019 $\pm$0.307
[dotted] 0.50 0.465 0.627 $\pm$0.251 0.174 -0.023 $\pm$0.196 0.300 -0.019 $\pm$0.307
[dotted] 0.55 0.462 0.657 $\pm$0.273 0.378 0.249 $\pm$0.307 0.281 -0.002 $\pm$0.292
[dotted] 0.60 0.489 0.698 $\pm$0.222 0.469 0.338 $\pm$0.371 0.415 0.137 $\pm$0.390
[dotted] 0.65 0.464 0.676 $\pm$0.218 0.465 0.411 $\pm$0.256 0.449 0.256 $\pm$0.357
[dotted] 0.70 0.458 0.657 $\pm$0.233 0.457 0.404 $\pm$0.212 0.369 0.297 $\pm$0.295
[dotted] 0.75 0.466 0.658 $\pm$0.197 0.541 0.550 $\pm$0.182 0.379 0.347 $\pm$0.319
[dotted] 0.80 0.429 0.591 $\pm$0.154 0.511 0.600 $\pm$0.196 0.444 0.551 $\pm$0.193
[dotted] 0.85 0.414 0.556 $\pm$0.201 0.468 0.558 $\pm$0.085 0.364 0.421 $\pm$0.270
[dotted] 0.90 0.295 0.504 $\pm$0.249 0.238 0.261 $\pm$0.070 0.316 0.401 $\pm$0.129
Average
[…]
Table 6: Pearson correlation coefficient between the CS (incl. different Match Scorers) and averaged human coverage scores for different matching thresholds, along with the evaluation runtime, on ArgKP21. For the scenario within topics and stances, standard deviations are indicated alongside the correlation values.
{tblr}
colspec = X[c,m,-1]|X[c,m,-1]|X[c,m]|X[c,m]|X[c,m], width=1stretch = 0, rowsep = 4pt, ArgSum System Scenario ARI Proportion Clustered Runtime
[…]
[r=3]m USKPM Excl. Noise 0.273 $\pm$0.045 0.884 $\pm$0.079 1.417 $\pm$0.136
[dotted] Incl. Noise 0.253 $\pm$0.057 1.000 $\pm$0.000 1.478 $\pm$0.113
[dotted] IC 0.268 $\pm$0.052 1.000 $\pm$0.000 4.377 $\pm$4.378
[…]
[dotted] Incl. Noise 0.547 $\pm$0.255 1.000 $\pm$0.000 1.133 $\pm$0.178
[…]
(BarH) Excl. Noise 0.725 $\pm$0.182 0.858 $\pm$0.064 53.86 $\pm$25.19
[dotted] Incl. Noise 0.640 $\pm$0.158 1.000 $\pm$0.000 53.86 $\pm$25.19
[…]
(SMatchToPr) Excl. Noise 0.682 $\pm$0.109 0.881 $\pm$0.042 2.723 $\pm$0.435
[dotted] Incl. Noise 0.639 $\pm$0.113 1.000 $\pm$0.000 2.723 $\pm$0.435
Table 7: Argument separation capability of clustering-based ArgSum systems for ArgKP21. While ARI refers to the highest ARI from several examined parameter settings, the proportion of clustered arguments and the clustering runtime refer to this highest ARI. All depicted values are averaged across topics and stances. Standard deviations are indicated behind each value.
{tblr}
colspec = X[c,m,-1]|X[c,m,-1]|X[c,m]|X[c,m]|X[c,m], width=1stretch = 0, rowsep = 4pt, ArgSum System Scenario ARI Proportion Clustered Runtime
[…]
[r=3]m USKPM Excl. Noise 0.298 $\pm$0.054 0.797 $\pm$0.051 2.090 $\pm$0.518
Incl. Noise 0.232 $\pm$0.049 1.000 $\pm$0.000 2.044 $\pm$0.507
IC 0.262 $\pm$0.054 1.000 $\pm$0.000 15.38 $\pm$9.633
[…]
Incl. Noise 0.315 $\pm$0.123 1.000 $\pm$0.000 1.931 $\pm$0.690
[…]
(BarH) Excl. Noise 0.411 $\pm$0.085 0.789 $\pm$0.079 201.9 $\pm$166.7
Incl. Noise 0.311 $\pm$0.099 1.000 $\pm$0.000 201.9 $\pm$166.7
[…]
(SMatchToPr) Excl. Noise 0.305 $\pm$0.071 0.783 $\pm$0.068 6.109 $\pm$2.323
Incl. Noise 0.245 $\pm$0.045 1.000 $\pm$0.000 6.109 $\pm$2.323
Table 8: Argument separation capability of clustering-based ArgSum systems for Debate. While ARI refers to the highest ARI from several examined parameter settings, the proportion of clustered arguments and the clustering runtime refer to this highest ARI. All depicted values are averaged across topics and stances. Standard deviations are indicated behind each value.
Figure 4: Highest ARI of clustering-based ArgSum systems from several examined parameter settings for ArgKP21. All depicted values are averaged across topics and stances. Standard deviations are indicated below the ARI values.
Figure 5: Highest ARI of clustering-based ArgSum systems from several examined parameter settings for Debate. All depicted values are averaged across topics and stances. Standard deviations are indicated below the ARI values.

### v2

@@
 B.3 LLM-based Evaluation
+For the evaluation, we only worked with user messages.
+User Message for Coverage Evaluation
+Your task is to evaluate a set of generated summaries obtained from a collection of arguments against a set of reference summaries.
+The evaluation is conducted according to the criteria of coverage, meaning that the set of generated summaries aims to cover the main statements contained in the set of reference summaries.
+Since each reference summary addresses a unique main statement, you are asked to count the number of reference summaries that are covered by the set of generated summaries.
+If a reference summary is only partially covered by the set of generated summaries, an increase of the count by 0.5 is allowed.
+Your counts aim to correlate well with human judgments.
+Make sure to always print the final count in the format "Coverage count: x.y" in a new line with no additional text in that line.
+Example:
+Set of Reference Summaries:
+  1. 1.
+    Banning guns would save lives
+  2. 2.
+    Guns can fall into the wrong hands
+  3. 3.
+    Guns lead to accidental deaths
+  4. 4.
+    Gun ownership allows for mass-shootings/general gun violence
+Set of Generated Summaries:
+  1. 1.
+    Banning guns would save thousands of lives
+  2. 2.
+    Some people do not know how to handle firearms. This is a danger to them and others.
+  3. 3.
+    Guns kill people, they should be banned
+  4. 4.
+    Firearms can fall into the hands of potential murderers
+  5. 5.
+    Firearms are a disgrace to humanity.
+  6. 6.
+    Without weapons, there would be no war.
+Coverage count: 3.5
+Evaluation Procedure:
+  1. 1.
+    Read the reference summaries. Do not print them again.
+  2. 2.
+    Read the generated summaries. Do not print them again.
+  3. 3.
+    Go through the set of reference summaries and determine whether the reference summary at hand is covered by at least one generated summary.
+  4. 4.
+    Once you have done this for each reference summary, count the number of covered reference summaries and return the resulting coverage count.
+Evaluation Task:
+Set of Reference Summaries:
+reference_summaries
+Set of Generated Summaries:
+candidate_summaries
+User Message for Redundancy Evaluation
+Your task is to evaluate a set of arguments on a certain debate topic and stance according to their uniqueness.
+Since arguments can be formulated differently, but address the same aspect of a debate, your task is to count the number of unique main statements addressed by the set of arguments.
+If a main statement addressed by an argument is only partially unique because it is also in parts covered by another argument, an increase of the count by 0.5 is allowed.
+Your counts aim to correlate well with human judgments.
+In the following, you are provided with an example, instructions for the evaluation procedure, and finally with your evaluation task.
+Example:
+Set of Arguments:
+  1. 1.
+    Banning guns would save lives
+  2. 2.
+    Guns can fall into the wrong hands
+  3. 3.
+    Guns lead to accidental deaths
+  4. 4.
+    Guns kill people, they should be banned
+  5. 5.
+    Gun ownership allows for mass-shootings/general gun violence
+  6. 6.
+    Some people do not know how to handle firearms. This is a danger to them and others.
+  7. 7.
+    Banning guns would save thousands of lives
+  8. 8.
+    Firearms can fall into the hands of potential murderers
+Number of Unique Main Statements: 4
+Explanation:
+  • Argument 1, 4, and 7 address the same main statement (guns kill people so without guns lives could be saved)
+  • Argument 2, 6, and 8 address the same main statement (guns could fall into the wrong hands, such as murders or people not knowing how to handle guns)
+  • Argument 3 addresses a unique main statement, focusing on accidents with guns
+  • Argument 5 addresses a unique main statement, focusing on intentional killing like terrorism or running amok
+Notes:
+  • Arguments 1, 4, and 7 are quite general, and therefore differ from the others
+  • E.g., argument 3 could also be assigned to 1, 4, and 7. Nevertheless, it focuses on accidents and is more specific
+Evaluation Procedure:
+  1. 1.
+    Read the arguments. Do not print them.
+  2. 2.
+    Go through the list of arguments, starting with the first argument.
+  3. 3.
+    Determine whether the argument at hand addresses a main statement of the debate.
+  4. 4.
+    Move on to the next one and consider whether it addresses a main statement and whether it has already been covered by previous arguments in the list.
+  5. 5.
+    Once you have done this for each argument, count the total number of unique main statements.
+  6. 6.
+    Return your uniqueness count in the format "Number of Unique Main Statements: x.y" in a new line with no additional text in that line.
+Always make this line the last line of your response and always include it.
+Evaluation Task:
+Set of Arguments: candidate_summaries
+Number of Unique Main Statements:
+Generation of Candidate Summaries and Reference Summaries
+Candidate Summaries and Reference Summaries are constructed by iterating over lists of generated and reference summaries, respectively.
+Each element in the list is formatted as an enumerated string, where each entry is prefixed with its index and a period.
+This ensures a structured representation of arguments or summaries for evaluation.
+Below is an example of how a list of three reference summaries would be converted into a formatted string:
+Set of Reference Summaries:
+  1. 1.
+    Renewable energy reduces carbon emissions.
+  2. 2.
+    Solar panels provide long-term cost savings.
+  3. 3.
+    Wind power is a reliable energy source.
+Similarly, a set of generated summaries follows the same structure, ensuring consistency in comparison.
 Appendix C Human Evaluation